can you put the following linear equation in slope intercept form:2x-5y=20
can you Solve for x in the following inequality: -17 < -2x + 3 < or equal to 25?

can you find the solution to what x is? 8 - x - 4 < 5

Answer by rfer(16322)

2x-5y=20
-5y=-2x+20
y=2/5x-4
------------
x > -1